Written Inc has outstanding 600,000 shares of $2 par common stock and 120,000 shares of no par 6% preferred stock with a stated value of $5. Dividends have been paid in every year except for the past two years and the current year.

Assuming that $270,000 will be distributed, and the preferred stock is cumulative and participating, how much will the common stockholders receive?

a) $162,000

b) $132,000 <-------------------------------- answer

c) $138,000

d) $72,000

Can someone help me with this question please? it says the answer is b) $ 132,000, but yet I'm getting $162,000????

below is my work :

600,000 x 6%= 36,000  

36,000 x 2 years = 72,000

270,000 - 72,000 - 36,000= $162,000

I KEEP getting $162,000 so I don't understand what I'm doing wrong. Can someone help please? show your work and thanks in advance!!!
Answer #1

preferred stock is cumulative and participating, You are considering the cumulative part in your solution but participating you are ignoring. In Participating  any dividends beyond the prescribed rate (6%) are shared on a ratio basis between preferred shareholders and common shareholders

Common Stock Holders = 1,200,000 x 6% = 72,000

270,000 - 72,000 - 36,000 - 72,000= $90,000 Remaining to be shared

Participation rate = 90,000/1,800,000 = 5%

1,200,000 par value x 5% = $60,000 owed to common shareholders

So Total Common Shareholders will receive = 72,000 + 60,000 = 132,000